OpenStudy (dls):

If f(x)=e^x + x and g is inverse of f then find g'(2+log2)

OpenStudy (dls):

\[\LARGE g'(2+\log2)=\frac{1}{e^{g(2+\log2)}+1}\] im stuck here..

OpenStudy (dls):

how do i get g(2+log2) ? when does f give 2+log2? :O

OpenStudy (experimentx):

x= - 2

OpenStudy (dls):

what happens at x=-2?

OpenStudy (dls):

f(-2)=1/e^2 -2?

OpenStudy (experimentx):

hold on a sec ... if by some manipulation i should be able to get that

OpenStudy (dls):

how can we get that value by simple hit and trial XD

OpenStudy (experimentx):

put x = log(2)

OpenStudy (dls):

woah :O

OpenStudy (dls):

biingo!